OpenStudy (anonymous):

2.1 meters x______=_______cm

OpenStudy (mathstudent55):

\( 2.1~m \times \dfrac{100~cm}{1~m} = ~~~?~~~cm \)

OpenStudy (anonymous):

yep I think thats it the way I wrote it is the way it looks on paper! confusing In the first blank would it be 100

OpenStudy (mathstudent55):

2.1 * 100 = ?

OpenStudy (anonymous):

thank you. you helped greatly